# Explorer: Scaling Exploration-driven Web Trajectory Synthesis for Multimodal Web Agents

### v2

@@ 5.1 In-domain Evaluation
As an intrinsic evaluation of the trajectory collection pipeline, we generate 100 test tasks using Explorer, disjoint from the train set. The SLM agents are tasked with executing the given tasks on live websites while an LLM-as-a-judge verifier (§ 3.2) evaluates the correctness of their actions at the trajectory level (Table 7). We observe that the fine-tuned agents significantly outperform their pre-trained counterparts. Thus, using in-domain web trajectory data training helps, which is a valuable sanity check. It also surpasses GPT-4o, further underscoring our synthetic data’s quality.
@@
Model Train Data Cross-Task Cross-Website Cross-Domain Avg.
Ele. Acc Op. F1 Step SR Ele. Acc Op. F1 Step SR Ele. Acc Op. F1 Step SR
In-Context Learning
GPT-3.5 $19.4$ $59.2$ $16.8$ $14.9$ $56.5$ $14.1$ $25.2$ $57.9$ $24.1$ $18.3$
GPT-4 $40.8$ $63.1$ $32.3$ $30.2$ $61.0$ $27.0$ $35.4$ $61.9$ $29.7$ $29.7$
SeeAct Zheng et al. (2024) $46.4$ $73.4$ $40.2$ $38$ $67.8$ $32.4$ $42.4$ $69.3$ $36.8$ $36.5$
Supervised Fine-Tuning
EDGE-9.6B Chen et al. (2024b) Syn. + M2W $30.0$ $21.1$ $22.4$ $24.5$
MiniCPM-GUI-3.1B Chen et al. (2024a) Syn. + M2W $23.8$ $86.8$ $20.8$ $20.3$ $81.7$ $17.3$ $17.9$ $74.5$ $14.6$ $17.6$
Scribe-Agent-L.-32B Shen et al. (2024) Syn. traj. $38.0$ $52.9$ $35.6$ $34.1$ $52.7$ $32.5$ $39.4$ $54.7$ $37.3$ $35.1$
AgentTrek-7B Xu et al. (2025) Syn. + M2W $60.8$ $88.9$ $55.7$ $57.6$ $88.1$ $51.4$ $56.0$ $87.5$ $52.6$ $53.2$
Explorer-4B M2W $48.1$ $88.0$ $44.8$ $49.1$ $87.2$ $45.0$ $46.9$ $87.7$ $44.6$ $44.8$
Explorer-4B Syn. + M2W $53.4$ $88.1$ $50.7$ $55.6$ $89.5$ $51.4$ $49.8$ $88.8$ $47.2$ $49.8$
Explorer-7B M2W $51.8$ $88.0$ $48.3$ $56.3$ $89.7$ $52.0$ $50.9$ $88.9$ $48.1$ $49.5$
Explorer-7B Syn. + M2W $56.5$ $90.3$ $53.2$ $60.5$ $90.7$ $56.7$ $55.7$ $90.4$ $53.0$ $54.3$
5.2 Mind2Web-Live Results
Improvement over base pre-trained models.
We observe that Explorer-7B yields improvements of 5.1% and 4.8% in average step success rate and key node completion rate, respectively, compared to the pre-trained Qwen2-VL-7B model. Similarly, Explorer-4B obtains gains of 15.5% and 15.9% in average step success rate and key node completion rate, respectively, over its pre-trained counterpart. In terms of full task success rate, Phi-3.5V improves significantly from 2.4% to 18.1%, while Qwen2-VL-7B improves from 14.5% to 19.3%. To the best of our knowledge, this represents the state-of-the-art performance on Mind2Web-Live for end-to-end web agents of this size that are trained exclusively on synthetic data.
Improvement over higher capacity pre-trained models.
Despite having much fewer parameters, we observe that Explorer-4B outperforms strong baselines such as Mistral-7B-Instruct-0.3 and Qwen2-72B-Instruct in full task SR by margins of 8.5% and 2.7%, respectively. The Phi-3.5V model obtains an 18.1% full task success rate, which is better than GPT-3.5 (15.4%), despite using orders of magnitude fewer parameters. The corresponding results for the entire set of 104 tasks, including unreachable websites, are given in Appendix A.
5.3 Multimodal-Mind2Web Results
Following Deng et al. (2023), we obtain the top-50 elements from a pre-trained DeBERTa He et al. (2021) candidate generation model which are then used to construct the accessibility tree and SoM image inputs. The results are shown in Table 9.
Among baselines, we include API-based models for in-context learning – GPT-3.5, GPT-4, and SeeAct Zheng et al. (2024). SeeAct is a web agent that performs web tasks using a two-step procedure of action generation and grounding using GPT-4V. Additionally, we include baselines that fine-tune small language models using synthetic data, followed by further fine-tuning on the Mind2Web training set. SeeClick Cheng et al. (2024) introduces a visual grounding model (Qwen-VL) trained on synthetically-generated grounding data. EDGE Chen et al. (2024b) synthesizes QA data on webpages to improve the grounded GUI understanding capabilities of MLLMs. ScribeAgent-Large Shen et al. (2024) and MiniCPM-GUI Chen et al. (2024a) use human-annotated trajectory data to train web agents. AgentTrek Xu et al. (2025) is a GUI agent baseline that also utilizes synthetic trajectory data to fine-tune SLMs for Mind2Web, similar to our setting. We observe that Explorer-7B fine-tuned on synthetic data from Explorer plus Mind2Web outperforms all baselines in average step success rate. Notably, it surpasses AgentTrek, which uses the same Qwen2-VL-7B MLLM backbone, highlighting the superior quality of our dataset. The broad domain coverage and task diversity in Explorer contribute to its superior generalization across environments.
@@ 5.4 MiniWob++ Results
Table 8 shows the results on the MiniWob++ benchmark in the zero-shot evaluation setting. Among baselines, we have API-based models, in-context learning using open-source LMs, and agentic models like AgentLM Zeng et al. (2024), CodeActAgent Wang et al. (2024b), Lemur-Chat Xu et al. (2024a) and AgentFlan Chen et al. (2024c) which include web-based demonstrations in their instruction tuning dataset. Synatra-CodeLlama-7B Ou et al. (2024) and AgentTrek Xu et al. (2025) also synthesize web-agent trajectories automatically. We observe that Explorer outperforms GPT-4 and general-purpose agents like AgentLM, CodeActAgent, Lemur-Chat, and AgentFlan. Explorer-4B surpasses Synatra-CodeLlama-7B and AgentTrek-7B despite using a much smaller model with 4.2B params, highlighting our synthetic data’s superior quality and potential for out-of-distribution (OOD) generalization.
Model Avg.  Step SR (%) Completion Rate (%) Full Task SR (%)
LLaVA-Mistral-7B $32.0$ $30.3$ $4.8$
Phi-3-mini (text-only) $36.6$ $34.0$ $13.3$
Phi-3.5V $44.0$ $39.4$ $18.1$
Qwen2-VL-7B $45.3$ $40.2$ $19.3$
Table 10: Ablation studies on language models used for fine-tuning (Mind2Web-Live).
@@ 6 Analyses
6.1 Ablation Studies
We conduct ablation studies to assess the impact of various design choices on overall performance (Table 10). To evaluate the importance of visual modality, we experiment with using just the textual modality for the Phi-3.5V model, replacing it with the text-only Phi-3-mini Abdin et al. (2024). In addition to Qwen2-VL-7B and Phi-3.5V, we also evaluate LLaVA-Mistral-7B Liu et al. (2023), a strong MLLM baseline. Our results show that omitting the visual modality leads to a sharp 4.8% drop in performance for Phi-3.5V, underscoring its importance for effective GUI grounding. Furthermore, LLaVA-Mistral-7B significantly underperforms compared to both Qwen2-VL-7B and Phi-3.5V, highlighting the necessity of a stronger MLLM backbone for improved GUI agent performance.
Figure 4: Experiments with data scaling using Explorer-4B on Mind2Web-Live. We experiment with using 100%, 50%, and 25% of the trajectory data. All results are averaged over three runs. All metrics exhibit improvement with increase in data scale.
6.2 Failure Modes of Trajectory Generation
We analyze cases where a generated trajectory is ultimately rejected by the task verifier agent. Our goal is to synthesize trajectory data that closely resembles human-annotated datasets for training web agents. However, since our pipeline collects trajectories through an exploration-driven approach, some trajectories result from random, incoherent action sequences that fail to align with a well-defined task intent. For instance, in shopping tasks, the agent may explore various products without demonstrating an intent to purchase (e.g., by adding items to the cart). Another failure case arises when the agent encounters errors on the final page due to automated browser detection, CAPTCHA verification, or an unresponsive website. We note that the verifier agent is instructed to judge a trajectory as successful if the task is completed, except for the final login and payment steps. The trajectories in failure modes are still valuable for web agents to learn low-level tasks such as form filling, basic interaction with web elements, and visual grounding.
Figure 5: Statistics for different error cases in Mind2Web-Live evaluation. Task deviation is the most prevalent error type.
6.3 Case Studies for Mind2Web-Live
Figure 5 presents the statistics for these error types.
@@ Acknowledgements
We would like to thank colleagues from the OSU NLP group and Microsoft for their feedback. This research was supported in part by ARL W911NF2220144 and computational resources from the Ohio Supercomputer Center Center (1987).
@@
This supplementary material provides additional details omitted in the main text.
@@
  • Appendix B: Trajectory Synthesis Cost Analysis
  • Appendix C: Reasoning Generation Agent
  • Appendix D: Prompt Details
@@ Appendix A Mind2Web Training and Evaluation Details
Mistral-7B-Instruct-0.3 Jiang et al. (2023) $33.0$ $28.6$ $25.0$ $11.5$
Qwen2-72B-Instruct Bai et al. (2023) $40.9$ $15.4$
@@ A.1 Mind2Web-Live
We exclude the following websites - https://www.kbb.com, https://www.sixflags.com, https://www.viator.com, https://www.menards.com, https://www.amctheatres.com, https://www.cargurus.com, https://www.gamestop.com, https://www.cabelas.com, https://www.rei.com due to denial of access faced during our tests. Table 6 shows the results on Mind2Web-Live for 83 out of 104 tasks across the remaining 37 websites. The results on the whole Mind2Web-Live evaluation set are given in Table A.1. The results in Table 6 are reported as the maximum over three runs, accounting for intermittent website access issues that may affect evaluation consistency. For Mind2Web-Live, the dataloader first samples training instances at the trajectory level and then randomly samples a step from the trajectory to construct the final training instance. Thus, the number of epochs is calculated at the trajectory level. We use a viewport resolution of $1280\times 720$ during inference. The Mind2Web-Live dataset is released under the MIT license which permits its use for academic research.
@@ A.2 Multimodal-Mind2Web
Following Deng et al. (2023), we obtain the top-50 elements from a pre-trained DeBERTa He et al. (2021) candidate generation model, which are then used to construct the accessibility tree and SOM image inputs. Following Ou et al. (2024), we always include the ground truth element in the input. We use a viewport resolution of $1280\times 720$ which includes the GT element during inference. We follow the setting in Zheng et al. (2024) and report element accuracy, operation F1, and step SR as evaluation metrics. All experiments on Multimodal-Mind2Web use a single training and evaluation run. The dataloader uniformly samples training instances from the set of action steps across all trajectories. The Multimodal-Mind2Web dataset is released under the Responsible AI license which permits its use for academic research.
@@
Appendix B Trajectory Synthesis Cost Analysis
@@
Appendix C Reasoning Generation Agent
Inspired by Xu et al. (2024b), the reasoning generation agent is a pre-trained Qwen2-VL-7B model that takes as input the current action, high-level task description, screenshot, accessibility tree, and action history. It then outputs a post-hoc reasoning trace for performing that action. These reasoning traces are helpful for training GUI agents in a chain-of-thought style.
Appendix D Prompt Details
The prompts for the task proposer agent, task refiner agent, task summarizer agent, and task verifier agent are given in Table D.5, Table D.7, Table D.8, and Table D.9, respectively. The training prompt for Explorer is given in Table D.10.
@@ Appendix D Prompt Details
1. Generate a single task that a user might perform on this webpage. Be creative and come up with diverse tasks
@@
2. The action should be an atomic action from the given vocabulary - click, type, select, scroll (up or down), or stop
@@
8. To input text, there is NO need to click the textbox first, directly type content. After typing, the system automatically hits the ‘ENTER‘ key.
@@
System Role You are an expert in evaluating the performance of a web navigation agent. The agent is designed to help a human user navigate a website to complete a task. Given the user’s intent, the agent’s action history, the final state of the webpage, and the agent’s response to the user, your goal is to decide whether the agent’s execution is successful or not.
@@
User Role The instruction is to {TASK DESCRIPTION}.
History actions: {PREVIOUS ACTIONS}
Here is the screen information: {ACCESSIBILITY TREE}

### v3

@@ 5.1 In-domain Evaluation
As an intrinsic evaluation of the trajectory collection pipeline, we generate $100$ test tasks using Explorer, disjoint from the train set. The SLM agents are tasked with executing the given tasks on live websites while an LLM-as-a-judge verifier (§ 3.2) evaluates the correctness of their actions at the trajectory level. Table 7 shows the results. We observe that the fine-tuned agents significantly outperform their pre-trained counterparts. Thus, using in-domain web trajectory data training helps, which is a valuable sanity check.
5.2 Mind2Web-Live Results
Improvement over base pre-trained models.
We observe that Explorer-7B yields improvements of $5.1$% and $4.8$% in average step success rate (SR) and key node completion rate, respectively, compared to the pre-trained Qwen2-VL-7B model. Similarly, Explorer-4B obtains gains of $15.5$% and $15.9$% in average step SR and key node completion rate, respectively, over its pre-trained counterpart. In terms of full task success rate, Phi-3.5V improves significantly from $2.4$% to $18.1$%, while Qwen2-VL-7B improves from $14.5$% to $19.3$%. To the best of our knowledge, this represents the state-of-the-art performance on Mind2Web-Live for models of this size trained exclusively on synthetic data.
@@ 5.2 Mind2Web-Live Results
Improvement over higher capacity pre-trained models.
Despite having much fewer parameters, we observe that Explorer-4B outperforms strong baselines such as Mistral-7B-Instruct-0.3 and Qwen2-72B-Instruct in full task SR by margins of $8.5$% and $2.7$%, respectively. The Phi-3.5V model obtains an $18.1$% full task success rate, which is better than GPT-3.5 ($15.4$%), despite using orders of magnitude fewer parameters. The corresponding results for the entire set of $104$ tasks, including unreachable websites, are given in Appendix A.1. We provide the ablation studies in Appendix A.3 and the error analysis in Appendix A.4.
Figure 3: Experiments with data scaling using Explorer-4B on Mind2Web-Live. We experiment with using $100\%$, $50\%$, and $25\%$ of the trajectory data. All results are averaged over three runs. All metrics exhibit improvement with an increase in data scale.
5.3 Multimodal-Mind2Web Results
Following Deng et al. (2023), we obtain the top-$50$ elements from a pre-trained DeBERTa He et al. (2021) candidate generation model, which are then used to construct the accessibility tree and SoM image inputs. The results are shown in Table 6.
Among baselines, we include API-based models for in-context learning – GPT-3.5, GPT-4, and SeeAct Zheng et al. (2024). SeeAct is a web agent that performs web tasks using a two-step procedure of action generation and grounding using GPT-4V. Additionally, we include baselines that fine-tune small language models using synthetic data, followed by further fine-tuning on the Mind2Web training set. SeeClick Cheng et al. (2024) introduces a visual grounding model (Qwen-VL) trained on synthetically-generated grounding data. EDGE Chen et al. (2024b) synthesizes QA data on webpages to improve the grounded GUI understanding capabilities of MLLMs. ScribeAgent-Large Shen et al. (2024) and MiniCPM-GUI Chen et al. (2024a) use human-annotated trajectory data to train web agents. AgentTrek Xu et al. (2025a) is a GUI agent baseline that also utilizes synthetic trajectory data to fine-tune SLMs for Mind2Web, similar to our setting. It synthesizes web trajectory data by guided replay from web tutorials. We observe that Explorer-7B fine-tuned on synthetic data from Explorer plus Mind2Web outperforms all baselines in average step success rate. Notably, it surpasses AgentTrek, which uses the same Qwen2-VL-7B MLLM backbone, highlighting the superior quality of our dataset. The broad domain coverage and task diversity in Explorer contribute to its superior generalization across environments.
@@ 5.3 Multimodal-Mind2Web Results
Statistic Count
# Unique task proposals $53$K
# Unique final task descriptions $81$K
# Final task descriptions $94$K
Table 9: Task diversity statistics at different stages of the data synthesis pipeline. Abstract task proposals evolve into diverse, fine-grained task descriptions as the agent explores the environment.
@@ 5.4 MiniWob++ Results
Table 8 shows the results on the MiniWob++ benchmark in the zero-shot evaluation setting. Among baselines, we have API-based models, in-context learning using open-source LMs, and agentic models like AgentLM Zeng et al. (2024), CodeActAgent Wang et al. (2024b), Lemur-Chat Xu et al. (2024) and AgentFlan Chen et al. (2024c) which include web-based demonstrations in their instruction tuning dataset. Synatra-CodeLlama-7B Ou et al. (2024) and AgentTrek Xu et al. (2025a) also synthesize web-agent trajectories automatically. We observe that Explorer outperforms GPT-4 and general-purpose agent baselines. Explorer-4B surpasses Synatra-CodeLlama-7B and AgentTrek-7B despite using a much smaller model with $4.2$B params, highlighting our synthetic data’s superior quality and strong potential for generalization to new web environments.
@@ 6 Analyses
Diversity Analysis.
We analyze the diversity of task descriptions across different stages of the data generation pipeline. At the initial proposal stage, the task proposer generates approximately $53$K unique high-level goals. As the agent explores the web environment, these proposals evolve into a total of $94$K final task descriptions, of which $81$K are unique. This progression shows that abstract task proposals can evolve into unique, fine-grained tasks with different constraints, shaped by the agent’s specific exploration path. Moreover, multiple trajectories corresponding to the same task description expose the agent to alternative solution paths, enhancing generalization.
Analysis of Verifier Accuracy.
Pred = Success Pred = Failure
GT = Success $0.39$ $0.05$
GT = Failure $0.14$ $0.42$
Table 10: Confusion matrix of the task verifier evaluated on a subset of $100$ Explorer-generated trajectories. We observe $81\%$ agreement with human judgment.
We did a human evaluation of the task verifier on a random set of $100$ trajectories generated using Explorer. We obtain $81\%$ agreement with human judgement for the task verifier, which is comparable to prior work Pan et al. (2024b); Xu et al. (2025a).
Failure Modes of Trajectory Generation.
To better understand failure cases in our pipeline, we perform a qualitative analysis of trajectories that were marked as unsuccessful by human annotators (Figure 4). The observed failure modes can be categorized as follows:
  • Grounding error during refinement: This error happens when the grounded action does not align with the natural language form of the agent’s action during the task refinement phase. This misalignment propagates to the summarizer, resulting in an inaccurate final task description.
  • Unresponsive website: The grounded action is correct and aligns with the task objective, but the website interaction fails due to non-responsive web elements during execution or dynamic content changes.
  • Summarization hallucination: The summarizer introduces extraneous constraints or goals to the task description that are not present in the underlying trajectory, causing misalignment.
  • Technical issues: The agent fails to complete the task due to external technical issues such as login requirements, media playback issues, or automated bot detection mechanisms.
  • Task incomplete: The agent reaches the step limit before completing all required actions, resulting in an incomplete trajectory.
Figure 4: Error type distribution in synthetic trajectory synthesis. Grounding errors in the refinement phase and summarization hallucinations are the most dominant error types.
@@ Acknowledgements
We would like to thank colleagues from the OSU NLP group and Microsoft Research for their valuable feedback. This research was supported in part by ARL W911NF2220144 and by computational resources provided by the Ohio Supercomputer Center Center (1987).
@@
In this supplementary material, we provide further details as follows:
@@
  • Appendix B: Cost Analysis
  • Appendix C: Task Complexity Analysis
  • Appendix D: System Prompts
@@ Appendix A Mind2Web Training and Evaluation Details
Figure A.1: Statistics for different error cases in Mind2Web-Live evaluation. Task deviation is the most prevalent error type.
@@
Mistral-7B-Instr. Jiang et al. (2023) $33.0$ $28.6$ $25.0$ $11.5$
Qwen2-72B-Instr. Bai et al. (2023) $40.9$ $15.4$
@@ A.1 Mind2Web-Live
We exclude the following websites - https://www.kbb.com, https://www.sixflags.com, https://www.viator.com, https://www.menards.com, https://www.amctheatres.com, https://www.cargurus.com, https://www.gamestop.com, https://www.cabelas.com, https://www.rei.com due to denial of access faced during our tests. Table 5 shows the results on Mind2Web-Live for $83$ out of $104$ tasks across the remaining $37$ websites. The results on the whole Mind2Web-Live evaluation set are given in Table A.1. The results in Table 5 are reported as the maximum over three runs, accounting for intermittent website access issues that may affect evaluation consistency. For Mind2Web-Live, the dataloader first samples training instances at the trajectory level and then randomly samples a step from the trajectory to construct the final training instance. Thus, the number of epochs is calculated at the trajectory level. We use a viewport resolution of $1280\times 720$ during inference. The Mind2Web-Live dataset is released under the MIT license, which permits its use in academic research.
@@ A.2 Multimodal-Mind2Web
Following Deng et al. (2023), we obtain the top-$50$ elements from a pre-trained DeBERTa He et al. (2021) candidate generation model, which are then used to construct the accessibility tree and SoM image inputs. Following Ou et al. (2024), we always include the ground truth element in the input. We use a viewport resolution of $1280\times 720$ which includes the GT element during inference. We follow the setting in Zheng et al. (2024) and report element accuracy, operation F1, and step SR as evaluation metrics. All experiments on Multimodal-Mind2Web use a single training and evaluation run. The dataloader uniformly samples training instances from the set of action steps across all trajectories. The Multimodal-Mind2Web dataset is released under the Responsible AI license, which permits its use in academic research.
@@
A.3 Ablation Studies
We conduct ablation studies to assess the impact of various design choices on overall performance (Table A.3). To evaluate the importance of visual modality, we experiment with using just the textual modality for the Phi-3.5V model, replacing it with the text-only Phi-3-mini Abdin et al. (2024). In addition to Qwen2-VL-7B and Phi-3.5V, we also evaluate LLaVA-Mistral-7B Liu et al. (2023), a strong MLLM baseline. Our results show that omitting the visual modality leads to a sharp $4.8$% drop in performance for Phi-3.5V, underscoring its importance for effective GUI grounding. Furthermore, LLaVA-Mistral-7B significantly underperforms compared to both Qwen2-VL-7B and Phi-3.5V, highlighting the necessity of a stronger MLLM backbone for better GUI agent performance.
Model Avg.  Step SR (%) Completion Rate (%) Full Task SR (%)
LLaVA-Mistral-7B $32.0$ $30.3$ $4.8$
Phi-3-mini (text-only) $36.6$ $34.0$ $13.3$
Phi-3.5V $44.0$ $39.4$ $18.1$
Qwen2-VL-7B $45.3$ $40.2$ $19.3$
Table A.3: Ablation studies on language models used for fine-tuning (Mind2Web-Live).
A.4 Case Studies for Mind2Web-Live
Figure A.1 presents the statistics for these error types.
Appendix B Cost Analysis
@@
Complexity level Count
Easy $8.2$K
Medium $44.3$K
Hard $41.2$K
Table B.5: Task complexity statistics. Most tasks fall within the medium to high complexity range.
Appendix C Task Complexity Analysis
Explorer contains web trajectories spanning multiple steps with an average of $7.7$ steps per trajectory (Table 3). Following prior work Zheng et al. (2024), we use the number of action steps in a trajectory as a proxy for task complexity, categorizing tasks as easy (2–4 steps), medium (5–7 steps), and hard (8–12 steps) in Table B.5. We observe that the majority of tasks fall within the medium to high difficulty range, indicating a high level of complexity in the dataset.
Appendix D System Prompts
The prompts for the task proposer agent, task refiner agent, task summarizer agent, task verifier agent, and captcha detection agent are given in Table D.7, Table D.9, Table D.10, Table D.11, and Table D.12, respectively. The training prompt for Explorer is given in Table D.13.
@@ Appendix D System Prompts
1. Generate a single task that a user might perform on this webpage. Be creative and come up with diverse tasks.
@@
System Role You are an expert in evaluating the performance of a web navigation agent. The agent is designed to help a human user navigate a website to complete a task. Given the user’s intent, the agent’s action history, and the final state of the webpage, your goal is to decide whether the agent’s execution is successful or not.
@@
User Role User Intent: {TASK_DESCRIPTION}
Action History: {ACTION_HISTORY}
The content of the last webpage in markdown format is given below: {LAST_PAGE_MARKDOWN}
The snapshots of all webpages corresponding to the actions are shown in the images: {SCREENSHOT_LIST}
@@
System Role You are an expert in evaluating whether the given webpage screenshot contains a captcha or not.
Given the last snapshot of the web page, your goal is to decide whether the webpage contains a captcha or not.
Output “Yes” if the given webpage shows a captcha, otherwise “No”.
IMPORTANT:
Format your response into a line as shown below:
Answer: “Yes” or “No”
User Role The screenshot of the web page is shown in the image.
{SCREENSHOT}
Table D.12: Prompt for Captcha Detection Agent.
@@
User Role The instruction is to {TASK_DESCRIPTION}.
History actions: {PREVIOUS_ACTION_LIST}
Here is the screen information: {A11Y_TREE}